Draw a supply and demand curve for wool suits in the country of Gauri. When international trade is allowed the price of a suit falls from 3 to 2 ounce of gold. Show a change in consumer surplus, producer surplus, and total surplus. How would a tariff on suit imports after these effects?
Answer #1

In following graph, AB & CD are domestic demand and domestic supply curves.

Autarky equilibrium is at point E where AB intersects CD, with pre-trade price P1 and quantity Q1.

Consumer surplus (CS) = Area between demand curve and pre-trade price = Area AEP1

Producer surplus (PS) = Area between supply curve and pre-trade price = Area CEP1

Total surplus (TS) = CS + PS = Area AEC

When free trade is allowed, new price is lower at P* corresponding to which domestic demand is Q2 and domestic supply is Q3, so imports = (Q2 - Q3).

New CS = Area between demand curve and world price = Area AFP* (Increase in CS = area P1EFP*)

New PS = Area between supply curve and world price = Area CGP* (Decrease in PS = area P1EGP*)

The tariff will increase the price of the good, and domestic price becomes Pt corresponding to which domestic demand is Q4 and domestic supply is Q5, so imports = (Q4 - Q5).

After tariff,

New CS = Area AHPt (Decrease in CS = Area PtHFP*)

New PS = Area CJPt (Increase in PS = Area PtJGP*)

Tariff revenue = Area PtHLP*

Deadweight loss from tariff = Area FHL + Area GJK